# 0.5 eV QCD Axion Cyclic Cosmology

### v46

@@ Abstract
A simple yet compelling physical picture is proposed for the nature of dark matter, dark energy, and the Big Bang. The proposal leads to a prediction with 4.4 parts per million precision for the 0.5 eV rest-mass energy of the quantum chromodynamic (QCD) axions that form the dark matter within this picture. A straightforward resolution of the Hubble tension follows from the agreement between the prediction for the present expansion rate of the universe within this picture and the best present determination from observations of the late universe.
@@ 1 Introduction
$H_{0}=71.4~506~(79)~{\rm km\cdot sec^{-1}\cdot Mpc^{-1}}~[{\rm 0.11~per~mille}].$ (1)
@@
Correspondingly, the predicted expansion rate in Eq. (1) differs sharply, by $8.8\sigma$, from the best present $\Lambda{\rm CDM}$ prediction using early-universe observations (Planck Collaboration et al., 2019, 2020)
@@
The straightforward resolution of the Hubble tension is that the late-universe observations are basically correct (on this point, see also D. Scolnic et al. (2025)), as are the $\Lambda{\rm CDM}$ predictions for the value of the cosmological constant and for the value of the present energy density of dark matter in the universe, but that the value of the present energy density of baryons is more than double the value predicted by $\Lambda{\rm CDM}$ (See Section 3 for the actual values of the corresponding cosmological parameters and for the comparison with $\Lambda{\rm CDM}$).
@@
where the cosmological parameter is $\Omega_{X}h^{2}=(8\pi/3)(Gu_{X}/c^{2})(h/H_{0})^{2}$ for energy density $u_{X}$, with $X=A,B,\gamma$, and it is $\Omega_{\Lambda}h^{2}=(c^{2}\Lambda/3)(h/H_{0})^{2}$ for the cosmological constant $\Lambda$. To resolve the Hubble tension, we propose a new cosmology, based on dark matter in the form of fundamental particles known as axions (S. Weinberg, 1978; F. Wilczek, 1978), that leads to predictions for the values of the cosmological parameters, $\Omega_{A}h^{2}$ and $\Omega_{B}h^{2}$, which have the same precision as the cosmological parameter for photons, $\Omega_{\gamma}h^{2}=2.473~(2)\times 10^{-5}~[{\rm 0.88~per~mille}]$, set by the precision of the best present estimate of the CMB temperature $T_{\gamma}=2.7255~(6)~{\rm K}~[{\rm 0.22~per~mille}]$ (D. J. Fixsen, 2009). Together with the first-principles prediction, within this axion cosmology, for the value of $\Omega_{\Lambda}h^{2}$, to an astounding precision of 44 parts per million (ppm), the prediction for $H_{0}$ given in Eq. (1) is obtained using Eq. (4).
@@
$\frac{\Omega_{B}h^{2}}{\Omega_{A}h^{2}}=\frac{2}{\pi^{2}}\left(x_{p}y_{T}\right)^{2}=0.393~565~531,$ (5)
where the numerically exact constant, $x_{p}=2.821~439~372$ (P. J. E. Peebles, 1993, pg. 159), gives the kinetic energy, $\check{E}=x_{p}kT_{\pi}$, for the axions on the peak of the black-body distribution at the temperature of the Big Bang, $T_{\pi}=y_{T}\left(\chi_{\rm QCD}\right)^{1/4}/k$, and the numerically exact constant, $y_{T}=(5/84)^{1/4}=0.493~938~274$, relates this temperature to the topological susceptibility of the quantum chromodynamic (QCD) vacuum, $\chi_{\rm QCD}=(m_{A}c^{2})^{2}f_{A}^{2}$ (S. Weinberg, 1978; F. Wilczek, 1978), given by the axion rest-mass energy, $m_{A}c^{2}$, and the axion decay constant, $f_{A}$. The value of this ratio within $\Lambda{\rm CDM}$ cosmology is roughly half as large (Planck Collaboration et al., 2019, 2020). This sharp difference resolves the Hubble tension, as shown in Eq. (1), and has implications for a wide range of astrophysical phenomena, including the formation of large-scale structure during the cross-over from radiation-domination to matter-domination (See Section 5), the synthesis of the nuclei of the chemical elements with low atomic mass during the Big Bang (See Section 6), and the spectrum of initial fluctuations in energy density at the Big Bang itself (See Section 7).
Figure 1: Near-infrared transmittance of polystyrene thin film at room-temperature and ambient pressure with no external static electric or magnetic fields (rough line) shows a decrease when the infrared light frequency is tuned to the resonant value $\nu_{A}\approx 4054~{\rm cm^{-1}}$ (vertical line, See Eq. (6)), for producing 0.5 eV QCD axion dark matter with the predicted rest-mass energy, $m_{A}c^{2}=hc\nu_{A}=0.502~580~(2)~{\rm eV}$ [4.4 ppm].
Yet, the most striking implication, perhaps, is the direct laboratory astrophysics observation of the production of dark matter (See Section 8). As near-infrared light passes through a material that has a static electric polarization within its structural unit cell, such as the common plastic material polystyrene, the magnetic field created by the light, ${\bf B}_{\rm NIR}$, couples to the electric field created by the polarization, ${\bf E}_{\rm DC}$, and their dot product, ${\bf E}_{\rm DC}\cdot{\bf B}_{\rm NIR}$, acts a coherent classical source for axions, provided that the energy of the photons in the light, $hc\nu_{A}$, with the wavenumber, $\nu_{A}$, matches the rest-mass energy, $m_{A}c^{2}$, of the axions. Indeed, a dip in the intensity of near-infrared light transmitted through a thin film of polystyrene was observed (Fig. 1) at roughly the predicted wavenumber
$\nu_{A}=4~053.~581~(18)~{\rm cm}^{-1}~[{\rm 4.4~ppm}],$ (6)
corresponding to the predicted rest-mass energy of the 0.5 eV QCD axion dark matter,
$m_{A}c^{2}=0.502~580~(2)~{\rm eV}~[{\rm 4.4~ppm}],$ (7)
and with roughly the size expected from the percent-level precise prediction for the strength of the axion-photon coupling (See Appendix D)
@@
In light of the apparent success of 0.5 eV QCD axion dark matter in resolving the Hubble tension, as well as the direct laboratory astrophysical evidence for the predicted axion rest-mass energy and axion-photon coupling, these astrophysical “constraints” and “exclusions” raise important questions about the conventional assumptions for a wide range of astrophysical phenomena. From the velocity dispersion of dark matter in the “halo” of the galaxy to the core helium abundance at formation for stars in globular clusters, these astrophysical assumptions underlie the analyses of the effects of axions on observations and simulations. Turning these analyses on their head, we conclude that the physical picture, presented in this Letter, stimulates a new look at these conventional assumptions, which cut across a wide swath of astrophysical phenomena from star formation to supernovae, from galactic structure to galactic chemical evolution, and from the structure to the evolution of pulsars, white dwarfs, red giants, and horizontal branch stars (See Section 10).
@@
Starting from the value $\check{E}$, given by Eq. (12), this kinetic energy reaches its peak value, $\hat{E}$, at future infinity when the dark matter dissolves and the universe begins contracting. For much of the past four decades, astrophysicists have looked to the Big Bang as a kind of cosmic accelerator for reaching collision energies sufficient to probe phenomena beyond the standard model of particle physics and the standard cold dark matter cosmology (L. Amendola et al., 2018; X. Chen et al., 2017; L. M. Lederman & D. N. Schramm, 1989, pg. 161). Yet, it is a simple historical fact that the last generation of the “classic” particle accelerators, known as synchrocyclotrons and built until the 1950s, already reached beam energies sufficient to make pairs of muons (L. M. Lederman & D. N. Schramm, 1989, pgs. 93–94), and, thus these classic accelerators already achieved seven decades ago collision energies greater than those typically experienced in the Big Bang, where the typical kinetic energy of radiation, $\check{E}$, was on the order of, but smaller than, the rest-mass energy of the muon $m_{\mu}c^{2}\approx 106~{\rm MeV}$ (See Eq. (12)).
@@
As expected, this value for the kinetic energy at the Big Bang in Eq. (16) is of the same order of magnitude as the estimate based on the value of the cosmological constant in Eq. (12). To go further, we use in Eq. (12) the actual values of the dimensionless numerically exact constants, $N_{U}=1/(2\pi^{2})=0.050~660~592$, and, $N_{n}=6\times 2!\zeta(3)/(\pi^{2}x_{p}^{3})=0.065~072~004$ (See Appendix A for the derivation of these values from the physical picture):
$\check{E}=104.~218~7~(8)~{\rm MeV}~\left(\frac{\Omega_{\Lambda}h^{2}}{0.35}\right)^{1/6}~[7.3~{\rm ppm}],$ (17)
@@
$=$ $0.117~542~(78)~[0.66~{\rm per~mille}],$
where the exact value of the astronomical unit ${\rm AU}=149~597~870.~700~{\rm km}$ (N. Capitaine et al., 2012) was used to derive the numerically exact value of the cosmological time, $h/H_{0}=({\rm AU}/100)(180/\pi)(60)^{2}\times 10^{6}~{\rm sec}=3.~085~677~581\times 10^{17}~{\rm sec}.$
As noted in Section 1, the high-precision prediction from first principles for the cold dark matter cosmological parameter in Eq. (19) agrees to within a percent or so with the best present estimate for this parameter within the standard $\Lambda{\rm CDM}$ cosmology using early-universe observations (Planck Collaboration et al., 2019, 2020)
@@
To push the precision of the first principles estimate of the cosmological constant and to determine the last remaining cosmological parameter, $\Omega_{B}h^{2}$, that of the baryons, we must leverage the numerically exact value for the baryon-to-axion energy density ratio, $\Omega_{B}h^{2}/\Omega_{A}h^{2}=0.393~565$, given in Eq. (5), and derived in Sec. 4:
@@
$=$ $0.046~260~(31)~[0.66~{\rm per~mille}],$
where the value predicted within our physical picture in Eq. (19) was used for $\Omega_{A}h^{2}$. As expected, this high-precision first principles prediction for the baryon abundance in Eq. (23) differs sharply, by more than $180\sigma$, roughly a factor of two, from the best present estimate for this parameter within the standard $\Lambda{\rm CDM}$ cosmology using early-universe observations (Planck Collaboration et al., 2019, 2020)
@@
To attain the promised 44 ppm precision for the value of $\Omega_{\Lambda}h^{2}$, it suffices to match the numerically exact baryon-to-axion energy density ratio, given in Eq. (5), with the value for this ratio computed using straightforward statistical mechanics in thermal equilibrium at the temperature of the Big Bang, $T_{\pi}=y_{T}(\chi_{\rm QCD})^{1/4}/k$, with the numerically exact dimensionless constant $y_{T}=(5/84)^{1/4}=0.493~938~274$ (See Appendix B for the derivation of $y_{T}$):
$\frac{\Omega_{B}h^{2}}{\Omega_{A}h^{2}}=\frac{1}{6}\left(\frac{m_{p}}{m_{A}}e^{-x_{p}}I(x_{p})+\frac{m_{n}}{m_{A}}e^{-x_{n}}I(x_{n})\right),$ (25)
where the axion-to-photon number density ratio, $n_{A}/n_{\gamma}=6$, has been used (See Appendix A for derivation), and the nucleon mass ratio, $x_{N}=m_{N}c^{2}/(kT_{\pi})$, for N=p,nN=p,~n, determines the equilibrium nucleon-to-photon number density ratio at the Big Bang, $e^{-x_{N}}I(x_{N})$, in terms of the numerically exact integral,
$I(x_{N})=\frac{1}{2!\zeta(3)}\int_{0}^{\infty}dxe^{-x}(x+x_{N})x^{1/2}(x+2x_{N})^{1/2}.$ (26)
Using the value of the rest-mass of the 0.5 eV QCD axion from Eq. (7), derived in Appendix B, the matching of the baryon-to-axion energy density ratio in Eq. (25) with the numerically exact value predicted in Eq. (5) then determines the temperature of the Big Bang, $T_{\pi}$, to 3.2 ppm, and the topological susceptibility of the QCD vacuum, $\chi_{\rm QCD}$, to 13 ppm:
$kT_{\pi}$ $=$ $36.~880~16~(12)~{\rm MeV}~[3.2~{\rm ppm}]$
$\sqrt{\chi_{\rm QCD}}$ $=$ $5~574.~947~(35)~{\rm MeV^{2}}~[6.3~{\rm ppm}],$ (27)
where the value of $\sqrt{\chi_{\rm QCD}}$ agrees with the best present estimate from chiral perturbation theory (CPT) $\sqrt{\chi_{\rm QCD}}({\rm CPT})=5~690~(50)~{\rm MeV}^{2}~[8.8~{\rm per~mille}]$ (M. Gorghetto & G. Villadoro, 2019), to within two percent, but is roughly 1400 times more precise. Repeating the quick and dirty calculation of the cosmological constant in Eq. (21), but now with the Big Bang kinetic energy, $\check{E}=x_{p}kT_{\pi}=104.~055~14~(33)~[3.2~{\rm ppm}]$, given by the first principles prediction for the Big Bang temperature in Eq. (3), yields the value of the cosmological parameter for the cosmological constant, $\Omega_{\Lambda}h^{2}$, to 44 ppm (See Appendix B for the derivation of the precision of the estimate):
$\Omega_{\Lambda}h^{2}$ $=$ $\left(0.35\right)\left(\frac{104.~055~14}{104.~218~7~(8)}\right)^{6}$ (28)
$=$ $0.346~717~(15)~[{\rm 44~ppm}].$
@@
$=$ $\frac{0.346~717}{(0.714~506~(79))^{2}}$
$=$ $0.679~146~(75)~[{\rm 0.22~per~mille}],$
where Eq. (28) was used for the value of $\Omega_{\Lambda}h^{2}$ and Eq. (1) was used for the value of $h$.
@@
To derive the numerically exact value of the baryon-to-axion energy density ratio, $\Omega_{B}h^{2}/\Omega_{A}h^{2}=(2/\pi^{2})(x_{p}y_{T})^{2}=0.393~565~531$, given in Eq. (5), which is the key to the proposed resolution of the Hubble Tension, presented in Sections 1, 2, and 3, we use the cyclic cosmology matching condition (See Appendix C):
$\left(\frac{T_{\pi}}{T_{\gamma}}\right)^{3}\left(\frac{\check{E}}{f_{A}}\right)^{2}\Omega_{B}h^{2}=4\left(\frac{\hat{E}}{f_{A}}\right)^{2}\Omega_{\Lambda}h^{2},$ (30)
@@
$=$ $\frac{2}{\pi^{2}}\left(\frac{\check{E}}{m_{A}c^{2}}\right)\left(\frac{\hat{E}^{2}}{E_{\rm Pl}^{2}}\right)$
[…]
$=$ $\frac{2}{\pi^{2}}\left(x_{p}y_{T}\right)^{2},$
@@
As the 0.5 eV QCD axion dark matter equation of state crosses over to matter-like at small redshifts from radiation-like at large redshifts, starting around $z_{\rm eq}$, as shown in Fig. 2, the dynamics of the expansion change to “matter-dominated” from “radiation-dominated.” Indeed, with six axions for each photon made in the cosmic background radiation (See Section 3 and Appendix A), axions dominate the expansion dynamics on both sides of “matter-radiation equality.” The cross-over in the axion equation of state (Fig. 2), beginning around $z_{\rm eq}$, drives the cross-over in the character of the expansion which then drives the large-scale formation seen by early-universe observations, just as the change from “radiation-domination” to “matter-domination” drives such structure to form in the standard cold dark matter cosmology (P. J. E. Peebles, 1993, pgs. 624–625).
@@
More than five decades ago, E. R. Harrison (1970) and Y. B. Zeldovich (1972) argued for a scale-invariant initial spectrum of fluctuations in the energy density of the radiation created in the Big Bang, $P(k)=Ak^{n_{s}}$, where the spectral index value, $n_{s}=1$, ensures scale-invariance, and the constant, $A$, is independnet of the co-moving wavevector of the fluctuation, $k$. Within the standard $\Lambda{\rm CDM}$ cosmology, the best present estimate from early-universe observations for the value of the spectral index, $n_{s}({\rm\Lambda CDM})=0.9677~(37)$ (Planck Collaboration et al., 2019, 2020), falls $8.8\sigma$ below the scale-invariant value, $n_{s}=1$. By contrast, the conformal symmetry of the cross-over from future infinity to the next big bang in 0.5 eV QCD axion cyclic cosmology (See Appendix C) requires the scale-invariant value, $n_{s}=1$, in agreement with the argument of E. R. Harrison (1970) and Y. B. Zeldovich (1972), but in sharp tension ($8.8\sigma$) with the standard $\Lambda{\rm CDM}$ value.
8 0.5 eV QCD Axion Laboratory Astrophysics
Figure 1 shows the laboratory astrophysics evidence for 0.5 eV QCD axion dark matter with the rest-mass energy, $m_{A}c^{2}$, predicted in Eq. (7), and the corresponding prediction for the axion-photon coupling strength, $g_{A\gamma\gamma}$, given in Eq. (8) and derived in Appendix D. The data were obtained with a standard commercial infrared spectrometer (Thermo-Fisher Scientific, Nicolet iS50R Fourier Transform Infrared Spectrometer, $0.125~{\rm cm^{-1}}$ resolution, $\rm KBr$} beamsplitter, DTGS photodetector at $293~{\rm K}$, Jacquinot stop diameter $w=1~{\rm mm}$), and a polystyrene thin film sample (International Crystal Laboratories, $t=38~{\mu m}$ thick, 293 K, 1 atm, in air, no external applied electric or magnetic fields, density $\rho=1.04~{\rm g/cm^{3}}$, unit cell molar mass $\mu=104~{\rm g/mol}$) that is similar to the standard reference material used to calibrate the infrared light frequency response of such spectrometers (NIST, 2025). The depth of the dip in transmittance, $\Delta T\approx 0.15~T$, relative to the background transmittance, $T\approx 82\%$, observed at roughly the value of the infrared light frequency predicted for the 0.5 eV QCD axion dark matter resonance, $\nu_{A}\approx 4054~{\rm cm}^{-1}$ (See Eq. (6)), with line-width, $\Delta\nu\approx 30~{\rm cm^{-1}}$, has roughly the expected size (F. Wilczek, 1978; S. Weinberg, 1978):
$\frac{\Delta T}{T}$ $=$ $\frac{1}{2}\left(m_{A}c^{2}g_{A\gamma\gamma}\right)^{2}\left(\epsilon_{0}E_{\rm DC}^{2}\right)\frac{NV}{hc\Delta\nu}\approx 0.15,$ (33)
where the static electric field within the unit cell of the plastic has strength $E_{\rm DC}\approx 5.7\times 10^{8}~{\rm V/cm}$ (J. L. Brédas & G. B. Street, 1985), the infrared beam within the plastic film has volume $V=(\pi/4)w^{2}t=3.0\times 10^{-5}~{\rm cm}^{3}$, and the number of unit cells within this volume is $N=N_{A}\rho V/\mu=1.8\times 10^{17}$.

### v47

# 0.5 eV QCD Axion Cosmology

@@ Abstract
A simple yet compelling physical picture is proposed for the nature of dark matter, dark energy, and the Big Bang. The proposal leads to a prediction with 0.53 parts per million precision for the 0.5 eV rest-mass energy of the quantum chromodynamic (QCD) axions that form the dark matter within this picture. A straightforward resolution of the Hubble tension within this picture follows from the agreement between the prediction for the present expansion rate of the universe and the best present determination from observations of the late universe.
@@ 1 Introduction
$H_{0}=71.931~(8)~{\rm km\cdot sec^{-1}\cdot Mpc^{-1}}~[{\rm 0.11~per~mille}].$ (1)
@@
Correspondingly, the predicted expansion rate in Eq. (1) differs sharply, by $10\sigma$, from the best present $\Lambda{\rm CDM}$ prediction using early-universe observations (Planck Collaboration et al., 2019, 2020)
@@
Within the physical picture presented in this Letter, the straightforward resolution of the Hubble tension is simply that the late-universe observations are basically correct (on this point, see also D. Scolnic et al. (2025)), as are the $\Lambda{\rm CDM}$ predictions for the value of the cosmological constant and for the value of the present energy density of dark matter in the universe, but that the value of the present energy density of baryons is more than double the value predicted by $\Lambda{\rm CDM}$ (See Section 3 for the actual values of the corresponding cosmological parameters and for the comparison with $\Lambda{\rm CDM}$).
@@
where the cosmological parameter is $\Omega_{X}h^{2}=(8\pi/3)(Gu_{X}/c^{2})(h/H_{0})^{2}$ for energy density $u_{X}$, with $X=A,B,\gamma$, and it is $\Omega_{\Lambda}h^{2}=(c^{2}\Lambda/3)(h/H_{0})^{2}$ for the cosmological constant $\Lambda$. To resolve the Hubble tension, we propose a new cosmology, based on dark matter in the form of fundamental particles known as axions (S. Weinberg, 1978; F. Wilczek, 1978), that leads to predictions for the values of the cosmological parameters, $\Omega_{A}h^{2}$ and $\Omega_{B}h^{2}$, which have the same precision as the cosmological parameter for photons, $\Omega_{\gamma}h^{2}=2.473~(2)\times 10^{-5}~[{\rm 0.88~per~mille}]$, set by the precision of the best present estimate of the CMB temperature $T_{\gamma}=2.725~5~(6)~{\rm K}~[{\rm 0.22~per~mille}]$ (D. J. Fixsen, 2009). Together with the first-principles prediction, within this axion cosmology, for the value of $\Omega_{\Lambda}h^{2}$, to an astounding precision of 22 parts per million (ppm), the prediction for $H_{0}$ given in Eq. (1) is obtained using Eq. (4).
@@
$\frac{\Omega_{B}h^{2}}{\Omega_{A}h^{2}}=\frac{2}{\pi^{2}}\left(x_{k}y_{T}\right)^{2}\left(\frac{1+\delta}{1+\epsilon}\right)=0.393~565~531~\left(\frac{1+\delta}{1+\epsilon}\right),$ (5)
where the small dimensionless quantity, $\epsilon=(4m_{p}/m_{\alpha}-1)Y_{P}\approx 0.002$, corrects for Big Bang nucleosynthesis of primordial helium with nucleus mass, $m_{\alpha}$, and primordial abundance by mass, $Y_{P}$, and the relativistic correction is $\delta\approx 0.06$ (See Appendix C), while the numerically exact constant, $x_{k}=2.821~439~372$ (P. J. E. Peebles, 1993, pg. 159), gives the kinetic energy, $\check{E}=x_{k}kT_{\pi}$, for the axions on the peak of the black-body distribution at the temperature of the Big Bang, $T_{\pi}=y_{T}\left(\chi_{\rm QCD}\right)^{1/4}/k$, and the numerically exact constant, $y_{T}=(5/84)^{1/4}=0.493~938~274$, relates this temperature to the topological susceptibility of the quantum chromodynamic (QCD) vacuum, $\chi_{\rm QCD}=(m_{A}c^{2})^{2}f_{A}^{2}$ (S. Weinberg, 1978; F. Wilczek, 1978), given by the axion rest-mass energy, $m_{A}c^{2}$, and the axion decay constant, $f_{A}$. The value of the corresponding ratio within $\Lambda{\rm CDM}$ cosmology is roughly half as large as the value given in Eq. (5) (Planck Collaboration et al., 2019, 2020). This sharp difference resolves the Hubble tension, as shown in Eq. (1), and has implications for a wide range of astrophysical phenomena, including the formation of large-scale structure during the cross-over from radiation-domination to matter-domination (See Section 5), the synthesis of the nuclei of the chemical elements with low atomic mass during the Big Bang (See Section 6), and the spectrum of initial fluctuations in energy density at the Big Bang itself (See Section 7).
A key prediction of the physical picture for dark matter, dark energy, and the Big Bang presented in this Letter is the value of the rest-mass energy of the 0.5 eV QCD axion dark matter to 0.53 parts per million (ppm) precision (See Appendix B for the derivation):
$m_{A}c^{2}=0.504~321~77~(27)~{\rm eV}~[{\rm 0.53~ppm}],$ (6)
This combines with the percent-level precise prediction for the strength of the axion-photon coupling (See Appendix D),
@@
Together these high-precision, first-principles predictions for the key physical parameters of axion dark matter provide, for the first time, the necessary inputs for the analysis of the astrophysical effects of axion dark matter.
@@
In light of the apparent success of 0.5 eV QCD axion dark matter in resolving the Hubble tension, these astrophysical “constraints” and “exclusions” raise important questions about the conventional assumptions for a wide range of astrophysical phenomena. From the velocity dispersion of dark matter in the “halo” of the galaxy to the core helium abundance at formation for stars in globular clusters, these astrophysical assumptions underlie the analyses of the effects of axions on observations and simulations. Turning these analyses on their head, we conclude that the physical picture, presented in this Letter, stimulates a new look at these conventional assumptions, which cut across a wide swath of astrophysical phenomena from star formation to supernovae, from galactic structure to galactic chemical evolution, and from the structure to the evolution of pulsars, white dwarfs, red giants, and horizontal branch stars (See Section 9).
@@
Starting from the value $\check{E}$, given by Eq. (11), this kinetic energy reaches its peak value, $\hat{E}$, at future infinity when the dark matter dissolves and the universe begins contracting. For much of the past four decades, astrophysicists have looked to the Big Bang as a kind of cosmic accelerator for reaching collision energies sufficient to probe phenomena beyond the standard model of particle physics and the standard cold dark matter cosmology (L. Amendola et al., 2018; X. Chen et al., 2017; L. M. Lederman & D. N. Schramm, 1989, pg. 161). Yet, it is a simple historical fact that the last generation of the “classic” particle accelerators, known as synchrocyclotrons and built until the 1950s, already reached beam energies sufficient to make pairs of muons (L. M. Lederman & D. N. Schramm, 1989, pgs. 93–94), and, thus these classic accelerators already achieved, seven decades ago, collision energies greater than those typically experienced in the Big Bang, where the typical kinetic energy of radiation, $\check{E}$, was on the order of, but smaller than, the rest-mass energy of the muon $m_{\mu}c^{2}\approx 106~{\rm MeV}$ (See Eq. (11)).
@@
As expected, this value for the kinetic energy at the Big Bang in Eq. (15) is of the same order of magnitude as the estimate based on the value of the cosmological constant in Eq. (11). To go further, we use in Eq. (11) the actual values of the dimensionless numerically exact constants, $N_{U}=1/(2\pi^{2})=0.050~660~592$, and, $N_{n}=6\times 2!\zeta(3)/(\pi^{2}x_{k}^{3})=0.065~072~004$ (See Appendix A for the derivation of these values from the physical picture):
$\check{E}=104.218~7~(8)~{\rm MeV}~\left(\frac{\Omega_{\Lambda}h^{2}}{0.35}\right)^{1/6}~[7.3~{\rm ppm}],$ (16)
@@
$=$ $0.117~949~(78)~[0.66~{\rm per~mille}],$
where the exact value of the astronomical unit ${\rm AU}=149~597~870.700~{\rm km}$ (N. Capitaine et al., 2012) was used to derive the numerically exact value of the cosmological time, $h/H_{0}=({\rm AU}/100)(180/\pi)(60)^{2}\times 10^{6}~{\rm sec}=3.085~677~581\times 10^{17}~{\rm sec}.$
As noted in Section 1, the high-precision prediction from first principles for the cold dark matter cosmological parameter in Eq. (18) agrees to better than a percent with the best present estimate for this parameter within the standard $\Lambda{\rm CDM}$ cosmology using early-universe observations (Planck Collaboration et al., 2019, 2020)
@@
To push the precision of the first principles estimate of the cosmological constant and to determine the last remaining cosmological parameter, $\Omega_{B}h^{2}$, that of the baryons, we must leverage the numerically exact value for the baryon-to-axion energy density ratio, $\Omega_{B}h^{2}/\Omega_{A}h^{2}=0.399~286~(14)$, given in Eq. (5), and derived in Sec. 4:
@@
$=$ $0.047~095~(31)~[0.66~{\rm per~mille}],$
where the value predicted within our physical picture in Eq. (18) was used for $\Omega_{A}h^{2}$. As expected, this high-precision first principles prediction for the baryon abundance in Eq. (22) differs sharply, by $190\sigma$, roughly a factor of two, from the best present estimate for this parameter within the standard $\Lambda{\rm CDM}$ cosmology using early-universe observations (Planck Collaboration et al., 2019, 2020)
@@
To attain the promised 22 ppm precision for the value of $\Omega_{\Lambda}h^{2}$, it suffices to match the numerically exact baryon-to-axion energy density ratio, given in Eq. (5), with the value for this ratio computed using straightforward statistical mechanics in thermal equilibrium at the temperature of the Big Bang, $T_{\pi}=y_{T}(\chi_{\rm QCD})^{1/4}/k$, with the numerically exact dimensionless constant $y_{T}=(5/84)^{1/4}=0.493~938~274$ (See Appendix B for the derivation of $y_{T}$):
$\frac{\Omega_{B}h^{2}}{\Omega_{A}h^{2}}=\frac{1}{6}\left(\frac{m_{p}}{m_{A}}e^{-x_{p}}J(x_{p})+\frac{m_{n}}{m_{A}}e^{-x_{n}}J(x_{n})\right)\frac{1+\delta}{1+\epsilon},$ (24)
where the axion-to-photon number density ratio, $n_{A}/n_{\gamma}=6$, has been used (See Appendix A for derivation), and the nucleon mass ratio, $x_{N}=m_{N}c^{2}/(kT_{\pi})$, for N=p,nN=p,~n, determines the ratio of the trace of the baryon stress-energy tensor to the photon number density at the Big Bang, $x_{N}e^{-x_{N}}J(x_{N})$, in terms of the numerically exact integral,
$J(x_{N})=\frac{1}{2!\zeta(3)}\int_{0}^{\infty}dxe^{-x}x^{1/2}(x+2x_{N})^{1/2}.$ (25)
Using the value of the rest-mass of the 0.5 eV QCD axion from Eq. (6), derived in Appendix B, the matching of the baryon-to-axion energy density ratio in Eq. (24) with the numerically exact value predicted in Eq. (5) then determines the temperature of the Big Bang, $T_{\pi}$, to 0.17 ppm, and the topological susceptibility of the QCD vacuum, $\chi_{\rm QCD}$, to 0.68 ppm:
$kT_{\pi}$ $=$ $36.911~503~1~(63)~{\rm MeV}~[0.17~{\rm ppm}]$
$\sqrt{\chi_{\rm QCD}}$ $=$ $5~584.420~4~(19)~{\rm MeV^{2}}~[0.34~{\rm ppm}],$ (26)
where the value of $\sqrt{\chi_{\rm QCD}}$ agrees with the best present estimate from chiral perturbation theory (CPT) $\sqrt{\chi_{\rm QCD}}({\rm CPT})=5~690~(50)~{\rm MeV}^{2}~[8.8~{\rm per~mille}]$ (M. Gorghetto & G. Villadoro, 2019), to within two percent, but is nearly 26000 times more precise. Repeating the quick and dirty calculation of the cosmological constant in Eq. (20), but now with the Big Bang kinetic energy, $\check{E}=x_{k}kT_{\pi}=104.143~568~(18)~[0.17~{\rm ppm}]$, given by the first principles prediction for the Big Bang temperature in Eq. (3), yields the value of the cosmological parameter for the cosmological constant, $\Omega_{\Lambda}h^{2}$, to 22 ppm (See Appendix B for the derivation of the precision of the estimate):
$\Omega_{\Lambda}h^{2}$ $=$ $\left(0.35\right)\left(\frac{104.143~568}{104.218~7}\right)^{6}$ (27)
$=$ $0.352~362~8~(78)~[{\rm 22~ppm}].$
@@
$=$ $\frac{0.352~362~8}{(0.719~31~(8))^{2}}$
$=$ $0.681~102~(75)~[{\rm 0.22~per~mille}],$
where we used Eq. (27) for $\Omega_{\Lambda}h^{2}$ and Eq. (1) for $h$.
@@
To derive the numerically exact value of the baryon-to-axion energy density ratio, $\Omega_{B}h^{2}/\Omega_{A}h^{2}$, given in Eq. (5), which is the key to the proposed resolution of the Hubble Tension, presented in Sections 1, 2, and 3, we use the cyclic cosmology matching condition (See Appendix C):
$\left(\frac{T_{\pi}}{T_{\gamma}}\right)^{3}\left(\frac{\check{E}}{f_{A}}\right)^{2}\left(\frac{1+\epsilon}{1+\delta}\right)\Omega_{B}h^{2}=4\left(\frac{\hat{E}}{f_{A}}\right)^{2}\Omega_{\Lambda}h^{2},$ (29)
@@
$=$ $\frac{2}{\pi^{2}}\left(\frac{\check{E}}{m_{A}c^{2}}\right)\left(\frac{\hat{E}^{2}}{E_{\rm Pl}^{2}}\right)\left(\frac{1+\delta}{1+\epsilon}\right)$
$=$ $\frac{2}{\pi^{2}}\left(\frac{\check{E}^{2}}{m_{A}c^{2}f_{A}}\frac{f_{A}}{\check{E}}\right)\left(\frac{f_{A}}{\hat{E}}\right)\left(\frac{1+\delta}{1+\epsilon}\right)$
$=$ $\frac{2}{\pi^{2}}\left(\frac{\check{E}}{\sqrt{\chi_{\rm QCD}}}\right)^{2}\left(\frac{f_{A}^{2}}{\check{E}\hat{E}}\right)\left(\frac{1+\delta}{1+\epsilon}\right)$
$=$ $\frac{2}{\pi^{2}}\left(x_{k}y_{T}\right)^{2}\left(\frac{1+\delta}{1+\epsilon}\right),$
@@
As the 0.5 eV QCD axion dark matter equation of state crosses over to matter-like at small redshifts from radiation-like at large redshifts, starting around $z_{\rm eq}$, as shown in Fig. 2, the dynamics of the expansion change to “matter-dominated” from “radiation-dominated.” Indeed, with six axions for each photon made in the cosmic background radiation (See Section 3 and Appendix A), axions dominate the expansion dynamics on both sides of “matter-radiation equality.” The cross-over in the axion equation of state (Fig. 2), beginning around $z_{\rm eq}$, drives the cross-over in the character of the expansion which then drives the large-scale structure formation seen by early-universe observations, just as the change from “radiation-domination” to “matter-domination” drives such structure to form in the standard cold dark matter cosmology (P. J. E. Peebles, 1993, pgs. 624–625).
@@
More than five decades ago, E. R. Harrison (1970) and Y. B. Zeldovich (1972) argued for a scale-invariant initial spectrum of fluctuations in the energy density of the radiation created in the Big Bang, $P(k)=Ak^{n_{s}}$, where the spectral index value, $n_{s}=1$, ensures scale-invariance, and the constant, $A$, is independent of the co-moving wavevector of the fluctuation, $k$. Within the standard $\Lambda{\rm CDM}$ cosmology, the best present estimate from early-universe observations for the value of the spectral index, $n_{s}({\rm\Lambda CDM})=0.9677~(37)$ (Planck Collaboration et al., 2019, 2020), falls $8.8\sigma$ below the scale-invariant value, $n_{s}=1$. By contrast, the conformal symmetry of the cross-over from future infinity to the next big bang in 0.5 eV QCD axion cyclic cosmology (See Appendix C) requires the scale-invariant value, $n_{s}=1$, in agreement with the argument of E. R. Harrison (1970) and Y. B. Zeldovich (1972), but in sharp tension ($8.8\sigma$) with the standard $\Lambda{\rm CDM}$ value.